# PromptPex: Automatic Test Generation for Language Model Prompts

### v1

@@ 2.4. Test Generation
Thus far, the prompt developer is able to align the intent with the input and output specifications extracted by PromptPex for the SOTA model. However, this does not account for how the prompt will actually perform on various inputs, especially when tested on other models. The test inputs generated by PromptPex gives the prompt developer a test suite designed to cover all the output constraints in the prompt and also to challenge the model to violate the constraints set within the prompt. By running these tests, the prompt developer can identify additional failures. Figure 5 shows a few tests generated by PromptPex for the POS prompt. This test suite also serves as a regression test suite for any future modifications to the prompt. With passing tests, the suite now encodes the developer’s intentions as concrete tests which were initially present only as specifications.
@@
These tests can now be executed across different models to compare and analyze the execution behavior of the prompt on other platforms. PromptPex assists the prompt developer in better understanding the behavior of the prompt by explicitly aligning the developer’s intentions with the extracted specification and implicitly through the generated test suite. This test suite is then used to identify differences between models and determine what modifications to the original prompt are needed to achieve the intended execution across various models as envisioned by the prompt developer.
@@ 2.5. Test Evaluation
One use case for PromptPex is to generate tests and then allow the user to add those tests to their existing test suite. Note that PromptPex does not generate the correct output for each test case, so this scenario requires the user to add the correct output for the generated tests as an additional step.
PromptPex supports an automated approach to test evaluation which checks test output not for correctness but for compliance with the prompt using an LLM. We discuss our approach to test evaluation for compliance in Section 3.
Figure 6. The end-to-end pipeline of PromptPex, which is implemented using a series of LLM prompts (shown in brown). The user provides their prompt (in green on the left), and PromptPex automatically generates input and output specifications, and tests (in blue), and can run the generated tests on multiple models (shown in purple).
3. Design
@@ 3.1. Input Specification
Similar to OR, the input specification (IS) describes the input and the constraints expected by the PUT. The IS defines what constitutes valid input. For the POS prompt, the IS, shown in Figure 3, describes the input as a sentence and a word, with the constraint that the word must be present in the sentence. Some prompts may also accept input in the form of a file; in such cases, the IS treats file input as string content and describes the content itself as the input rather than the file. PromptPex also generates the IS using gpt-4o. We frame it as a task to extract the IS to create valid inputs. We restrict any details about the output or how the input will be used in the computation inside the PUT. In creating the IS, we first extract a description of what the inputs are. If they are composed of multiple components, those components should be listed, and their constraints and properties should be described. The complete prompt for extracting the IS is available in Appendix A.2. Sometimes the PUT will attempt to handle corner cases. For example, in the POS prompt, if the word cannot be tagged with the listed tags, it returns Unknown. In this case, the IS extractor might incorrectly assume that words which cannot be tagged are not valid inputs. While this may be true, the prompt does not imply that. We explicitly added this case in the IS extractor prompt to consider such inputs as part of the input domain, even if there is a rule against them.
Both the OR and the IS are editable by the prompt developer, allowing them to modify the extracted specifications or augment them with constraints from the environment.
@@ 3.2. Output Specification or Rules
The output specification or output rules (OR) describe the constraints on the output generated for PUT. Rules are concrete, checkable, general, input-agnostic, and independent constraints over the output described in natural language and derived from the PUT. They are similar to assertions and post-conditions in a program, detailing what the output must be and must look like, without regard for how it is generated or derived. In our running example of the part-of-speech (POS) prompt, the OR, as shown in Figure  4, includes rules stating that the output should either be a POS tag, CantAnswer, or Unknown. It is important to note that OR do not capture how the output is generated; instead, it focuses solely on constraints over the generated output. This quality of OR helps in keeping it separate from the input, allowing evaluation of the output based solely on the OR, regardless of input. The groundedness of the rules within the OR in the prompt is a necessary condition for the OR itself as it implies that all rules are valid and are present in the prompt, while the exhaustiveness in accurately covering all rules from the prompt is the sufficient condition for the OR as no more rules are required.
We use gpt-4o for extraction of the OR from PUT. The complete prompt is available in Appendix A.4. We frame it as a task to extract the rules for output validation such that the inputs are not available. We enforce the following properties during the extraction of the rules:
  • If an example is present in the prompt, do not generate rules specifically for that example. Generalize them so that they will applicable for other possible inputs.
  • Rules must be clear, concrete and independent from each other such that they can individually be used to validate the output.
  • They should not contain any information about how the output depends on the input or how the output is computed.  [^4] We limit our scope to output compliance testing where we only validate the output unlike functional testing where we validate the output for the given input.
@@ 3.3. Test Generation
Tests for PUT are generated by PromptPex using OR and IS. The OR is utilized to create directed tests that challenge the model to correctly adhere to each rule, while the IS is used to create valid test cases. The input domain specified by the IS is usually extensive, as these inputs are in natural language, which can be presented in multiple forms. It is also non-trivial to determine which part of the prompt a given input covers, leaving us without a notion of coverage in prompts unlike what we have in programs.
@@ 3.3.1. Exhaustiveness
Given the vast range of possible inputs, creating an exhaustive set of test cases is challenging. To achieve this, we generate tests for each rule in the OR. We argue that if our OR is exhaustive (completely covers the prompt), the tests generated for the rules in the OR are also exhaustive.
PromptPex not only generates tests but also associates each test with a specific rule, providing reasoning for its creation. Beyond exhaustiveness, this approach allows for various analyses of the tests, as they are directly linked to a rule. We use this approach to attempt to develop an exhaustive test suite. Other possible future uses include modular updates to the test suites, allowing the addition of new tests for new rules and the removal of old tests linked to obsolete rules.
@@ 3.3.2. Generating challenging tests
Figure 7. Example illustrating the use of inverse rules for test generation. Because our inverse rule was used to generate the test in the lower half of the figure, the test intentionally focuses on the word ”Play” which depending on context can be labeled with multiple parts of speech.
Inverse Rules: We generate inverse rules from the given rules. The inverse of a rule is a semantic inversion that violates the rule by describing its opposite. For example, if a rule in OR states that the output must always be a tag, it can be inverted into a rule that enforces the output to be the tag along with the actual name of the part of speech. PromptPex uses gpt-4o for generating inverses of the given rules. We ask it to generate inverse rules such that they contradict the given rules. The complete prompt is available in the Appendix A.5.
We generate tests for both the rules from the OR and their inverses as shown in Figure 7. This approach helps us generate tests that cover intriguing test cases which might cause a given model to violate the OR.
@@ 3.3.3. Valid Tests
A test may not follow the IS and might create output that violates the rules described in the OR and PUT. Such test cases are not the most valuable, as the prompt developer knows that some of those inputs can be pre-filtered before they are passed to the prompt. However, these test cases might be useful for testing any input validation pipeline that processes the input or for identifying gaps in the assumed input domain. The most valuable tests are those that follow the IS, as they are directly actionable and require fixing. We use the extracted IS to guide PromptPex to generate tests that strictly follow the IS. Even in scenarios where the prompt-based system is exposed to raw inputs, the same IS can validate inputs and disregard any input that does not meet the input specification.
@@
The test generator in PromptPex uses gpt-4o and takes a rule (which can also be an inverse rule), IS and PUT as input. It generates a test along with the reasoning of why the model thinks this particular test during execution will comply with the rule it was generated for. We ask the test generator to start with first understanding: What is an input? What are the different components of a valid input? What are the syntax and semantics related constraints from IS? In the prompt, we direct the model to consider these issues during test generation. Each test must be generated such that the expected output or behavior demonstrates adherence to the given rules for a range of scenarios, including boundary cases, typical cases, and edge cases. The exact prompt used is available in the Appendix A.5.
3.4. Test Evaluation: Compliance versus Correctness
The generated test cases can then be executed on different models. Running tests on different models helps the prompt developer understand the behavior of diverse test inputs to the prompt when run on multiple models.
@@ 3.4. Test Evaluation: Compliance versus Correctness
As mentioned, because we lack the correct output for each test input, our automated strategy for test evaluation is an LLM-based determination of test compliance. Compliance is determined by our LLM-as-a-judge (llm-as-a-judge, ) which, given the original prompt and the test output, determines whether the output complies with the requirements described in the prompt. As a result, our evaluation explores only a partial understanding of the model output and must be augmented with additional testing. However, as demonstrated in traditional software engineering, testing the assertions over the output can also be very valuable (assertions-swe-burnett2003end, ; assertions-swe-clarke2006historical, ; assertions-swe-taromirad2024literature, ). After the tests are run, the results of our validator can assist the prompt developer in updating the prompt to correctly handle tests with invalid results.
@@
To implement LLM-as-a-judge, we use gpt-4o and provide it the output generated by executing a test and the prompt (PUT) for which the test was executed. To allow a fair comparison with a test generator that does use PromptPex, we ensured that no artifacts specific to PromptPex are used in the validation and that the test output validator is generic, applicable to any prompt and test output regardless of the input, the method used to generate the test, or the model used to run the test.
This validator is used to evaluate the outputs of tests generated by both the baseline and PromptPex. We enforce the following properties for this output validator:
  • Keep the evaluation independent of the input, as it will not be provided.
  • Must not speculate, infer, or make any assumptions during the evaluation.
  • Always check for compliance, not correctness.
The complete prompt can be found in Appendix A.1. The output validator generates a boolean result (compliant/non-compliant) and explanation for the decision.
@@ 4.1. Evaluation Goals
  • RQ1: Does having specifications help in generating better tests?
  • RQ2: How useful are rules for generating better tests?
  • RQ3: Does having an input specification and using it to generate tests create more valid tests?
  • RQ4: Can automatically generated tests help determine if a model is suitable for a prompt?
4.2. Baseline
In our evaluation, we used a zero-shot LLM-based test generator as the baseline. We use gpt-4o to generate tests for both the baseline and PromptPex. We instruct the model to develop multiple test cases by inferring the functional specification and input for the given prompt. We ask it to begin with understanding possible inputs and its components, what the syntax and semantics related constraints are, and possible input scenarios. We enforce the following properties while generating the baseline tests:
  • The test cases must be designed to validate whether the output properly adheres to description.
  • A good test must always be a valid input meeting the requirements mentioned in the description.
  • The test cases must be diverse and distinct.
  • Each test case must be crafted to rigorously assess whether the output meets the stipulated behavior based on the provided prompt.
  • The input scenarios used for creating the tests must be valid, realistic, and fully comply with the given description.
  • Generate test cases to broadly cover a range of scenarios, including boundary cases, typical cases, and edge cases, to thoroughly evaluate the software’s adherence to the description under various conditions.
  • Each test case should adhere to principles of good software testing practices, emphasizing coverage, specificity, and independence.
  • Focus on creating diverse test cases that effectively challenge the prompt’s capabilities by critically assessing potential weaknesses in the handling of inputs by the prompt.
The complete prompt is available in Appendix C.1. We chose this as our baseline because LLMs are capable of interpreting the prompt and generating test cases for them. LLMs are already found useful in generating unit tests for traditional software (llm-old-swe-test-gen-og, ; llm-old-swe-test-gen-siddiq2024using, ; llm-old-swe-test-gen-tang2024chatgpt, ; llm-old-swe-test-gen-xie2023chatunitest, ; llm-old-swe-test-gen-yuan2023no, ; llm-old-swe-test-gen-yuan2024evaluating, ; llm-old-swe-test-gen-yang2024evaluation, ). We ensured that the prompt used to generate the baseline tests is robust and underwent multiple revisions to enhance its ability to generate effective tests. To provide a fair comparison with PromptPex, we explicitly covered all requirements that are implicitly enforced by PromptPex, such as generating valid, challenging, and diverse test cases that comprehensively cover the prompt and help identify any flaws. Although there will always be opportunities to improve the baseline prompt, the same holds for the prompts used in PromptPex. We have refined both approaches sufficiently so that they can be effectively compared.
4.3. Metrics
@@
We define effective tests as those that expose limitations in the prompt, which means that they result in more failures. We consider non-compliance with the prompt as the metric for test quality. This includes any violations of the rules and constraints described in the prompt. This approach is similar to checking all the assertions over the output generated by a traditional program.
@@
We used LLM-as-a-judge to evaluate whether the output generated by the prompt violates any rules or constraints within the prompt as described in Section 3.4. We could have also developed a validator that does not use the PUT but instead relies on the OR, as they precisely represent the constraints over the output. However, since the OR is part of PromptPex and was not provided to the baseline test generator, we refrained from using it to maintain a fair and equal comparison.
We run the tests generated by PromptPex and the baseline on gpt-4o-mini, gemma2:9b, qwen2.5:3b, llama3.2:1b. We compare the non-compliance of the output from the generated tests. In our evaluation, we consider more test non-compliance as better since it indicates that the tests are more effective at identifying gaps in the prompt’s constraints.
4.3.2. Test Validity
Valid tests follows the input specification defined in the PUT. Valid tests are important as they provide instant feedback about the PUT and needs to be fixed. We also consider the test validity as a metric for test quality of the tests generated by PromptPex.
We use the IS with an LLM-as-a-judge to determine if the generated test is a valid input. The input validation prompt is available in Appendix C.2. We specifically used IS as PromptPex allows editing of generated IS to capture the precise input specification including implicit constraints arising from preprocessed inputs.
4.3.3. Groundedness and Spec Agreement
In addition to measuring test non-compliance and test validity which are direct indicators of the quality of the tests generated, we also evaluate the generated OR as these are directly used to generate tests by PromptPex and have impact on the quality of the generated tests.
@@ 4.3.3. Groundedness and Spec Agreement
A rule in the OR is considered grounded if it is present in the PUT. We do not want to generate tests for the rules which are not present in the PUT. We consider a higher groundedness of the rules as an indirect metric for test quality. To determine if the generated rules are grounded in the original PUT, we used LLM-as-a-judge and ask it to confirm rule groundedness. The complete prompt for checking the groundedness of the rules is available in Appendix  C.3.
@@
Figure 8. Overview of the spec agreement estimation for the tests generated by PromptPex. The specifications are extracted from the prompt and a spec prompt is created by appending the output specification to the task specification. The valid baseline tests are then executed with the original and synthetic prompt. Spec agreement is considered high when the spec prompt and the original prompt behave similarly. This is estimated by comparing the number of non-compliant test results.
@@
Another goal for generating OR is to ensure that all important constraints mentioned in the PUT are captured in the OR. We refer to this as spec agreement. To compute this, we extract a description of how the output must be computed from the PUT, the task specification. The prompt for extracting the task specification is available in Appendix C.4. We append the task specification with our extracted OR to derive a spec prompt which in theory should capture all the same constraints as the PUT. We do not use the IS in the spec prompt because PromptPex generates tests for each rule in the OR. To estimate how closely the extracted specification represents the original PUT, we compare the spec prompt with the original PUT. We feed tests generated from our baseline test generator to both the spec prompt and the PUT to compare their behaviors as shown in Figure 8. We use the baseline tests instead of the PromptPex tests so that they are not influenced by our process of generating them from the OR. The spec prompt and the original PUT must have similar behavior for the generated OR to have high spec agreement. We use cosine similarity of the non-compliance percentage of spec prompt and the original PUT to derive a spec agreement score.
Name of Prompt Prompt Description Source
speech-tag Determine the part of speech for a given word within a sentence using a predefined set of tags. The task may return tags such as noun, verb, adjective among others, or return ”Unknown” or ”CantAnswer” if the word cannot be classified. Modified from an example used by Schnabel et. al. (prompt-opt-sammo-schnabel2024symbolic, )
text-to-p Format a paragraph of text into HTML by splitting it into sentences and wrapping each sentence with paragraph tags, while enhancing key words and phrases with additional HTML tags. format_text prompt from GhostWriter (arxiv2024ghostwriter, )
shakespeare Assist users in creating text that mimics the Shakespearean style of writing, including the use of archaic language and stylistic elements typical of the period. Azure AI Studio Prompt Catalog (azure2023shakespeare, )
sentence Rewrite a sentence to improve its readability and make it more conversational while preserving its original meaning. This includes simplifying complex phrases and enhancing engagement through fluid structure. The Big Prompt Library (bigpromptlibrary2023sentence, )
extract-names Extract model names from machine learning paper abstracts, returning a structured array of identified model names or ”NA” if none are found. Information extraction prompt from Prompt Hub (promptguide2023extract, )
elements Extract important entities from a text, including company names, people names, specific topics, and general themes, presented in a structured list format. OpenAI documentation (openai2023bestpractices, )
classify Classify a news article into one of several predefined categories such as World, Sports, Business, or Sci/Tech, based on its content and context. Prompt used in a tutorial (pupuweb2023classification, )
art-prompt Create detailed prompts based on user descriptions for generating AI images, focusing on key characteristics, timing, lighting, and the desired emotional impact of the image in a concise single paragraph. The Big Prompt Library (bigpromptlibrary2023artprompt, )
Table 1. Description of the prompts used in the evaluation with their sources.
4.4. Benchmarks

### v2

@@ 2.4. Test Generation
So far, the prompt developer is able to align the intent with the input and output specifications extracted by PromptPex for the SOTA model. However, this does not account for how the prompt will actually perform on various inputs, especially when tested on other models. The test inputs generated by PromptPex gives the prompt developer a test suite designed to cover all the output constraints in the prompt and also to challenge the model to violate the constraints set within the prompt. By running these tests, the prompt developer can identify additional failures. Figure 5 shows a few tests generated by PromptPex for the POS prompt. This test suite also serves as a regression test suite for any future modifications to the prompt. With passing tests, the suite now encodes the developer’s intentions as tests which were initially present only as specifications.
@@
These tests can now be executed across different models to compare and analyze the execution behavior of the prompt on other platforms. PromptPex assists the prompt developer in better understanding the behavior of the prompt by explicitly aligning the developer’s intentions with the extracted specification and implicitly through the generated test suite.
@@ 2.5. Test Evaluation
PromptPex supports an automated approach to test evaluation which checks test output not for correctness but for compliance with the prompt using an LLM. One use case for PromptPex is to generate tests and then allow the user to add those tests to their existing test suite. Note that PromptPex does not generate the correct output for each test case, so this scenario requires the user to add the correct output for the generated tests as an additional step.
Figure 6. The end-to-end PromptPex pipeline which is implemented using a series of LLM prompts (shown in brown). The user provides their prompt (in green on the left), and PromptPex automatically generates input and output specifications, and tests (in blue), and can run the generated tests on multiple models (shown in purple).
3. PromptPex Design
@@ 3.1. Input Specification
The IS describes the input and the constraints expected by the PUT. The IS defines what constitutes valid input. For the POS prompt, the IS, shown in Figure 3, describes the input as a sentence and a word, with the constraint that the word must be present in the sentence. Some prompts may also accept input in the form of a file; in such cases, the IS treats file input as string content and describes the content itself as the input rather than the file. PromptPex generates the IS using gpt-5_2025-08-07. We frame it as a task to extract the IS to create valid inputs. We restrict any details about the output or how the input will be used in the computation inside the PUT. In creating the IS, we first extract a description of what the inputs are. If they are composed of multiple components, those components should be listed, and their constraints and properties should be described. Sometimes the PUT will attempt to handle corner cases. For example, in the POS prompt, if the word cannot be tagged with the listed tags, it returns Unknown. In this case, the IS extractor might incorrectly assume that words which cannot be tagged are not valid inputs. While this may be true, the prompt does not imply that. We explicitly added this case in the IS extractor prompt to consider such inputs as part of the input domain, even if there is a rule against them. Both OR and IS can be edited by the prompt developer, allowing them to modify the extracted specifications or augment them with constraints from the environment.
@@ 3.2. Output Specification or Rules
OR describes the output constraints generated for PUT. Rules are concrete, checkable, general, input-agnostic, and independent constraints over the output described in natural language and derived from the PUT. They are similar to assertions and post-conditions in a program, detailing what the output must be, without regard for how it is generated or derived. In our POS prompt, the OR, as shown in Figure  4, includes rules stating that the output should either be a POS tag, CantAnswer, or Unknown. It is important to note that OR do not capture how the output is generated; instead, it focuses solely on constraints over the generated output. This quality of OR helps in keeping it separate from the input, allowing evaluation of the output based solely on the OR, regardless of input. The groundedness of the rules within the OR in the prompt is a necessary condition for the OR itself as it implies that all rules are valid and are present in the prompt, while the exhaustiveness in accurately covering all rules from the prompt is the sufficient condition for the OR as no more rules are required.
We use gpt-5_2025-08-07 for extraction of the OR from PUT. We frame it as a task to extract the rules for output validation such that the inputs are not available. We enforce the following properties during the extraction of the rules: 1) If an example is present in the prompt, do not generate rules specifically for that example. Generalize them so that they will applicable for other possible inputs, 2) Rules must be clear, concrete and independent from each other such that they can individually be used to validate the output. and 3) They should not contain any information about how the output depends on the input or how the output is computed.  [^3] We limit our scope to output compliance testing where we only validate the output unlike functional testing where we validate the output for the given input.
@@ 3.3. Test Generation
Tests for PUT are generated by PromptPex using OR and IS. The OR is utilized to create directed tests that challenge the model to adhere correctly to each rule, while the IS is used to create valid test cases. The input domain specified by IS is usually extensive, as these inputs are in natural language, which can be presented in multiple forms. It is also non-trivial to determine which part of the prompt a given input covers, leaving us without a notion of coverage in prompts unlike what we have in programs.
@@ 3.3.1. Exhaustiveness
Given the vast range of possible inputs, creating an exhaustive set of test cases is challenging. To achieve this, we generate tests for each rule in OR. We argue that if our OR is exhaustive (completely covers the prompt), the tests generated for the rules in the OR are also exhaustive. PromptPex not only generates tests but also associates each test with a specific rule, providing reasoning for its creation. Beyond exhaustiveness, this approach allows for various analyses of the tests, as they are directly linked to a rule. We use this approach to attempt to develop an exhaustive test suite. Other possible future uses include modular updates to the test suites, allowing the addition of new tests for new rules and the removal of old tests linked to obsolete rules.
@@ 3.3.2. Generating challenging tests
Figure 7. Example illustrates inverse rules for test generation. Because inverse rule was used to generate the test in the lower half of the figure, the test intentionally focuses on the word ”Play” which can be labeled with multiple parts of speech.
Inverse Rules: We generate inverse rules from the given rules. The inverse of a rule is a semantic inversion that violates the rule by describing its opposite. For example, if a rule in OR states that the output must always be a tag, it can be inverted into a rule that enforces the output to be the tag along with the actual name of the part of speech. PromptPex uses gpt-5_2025-08-07 for generating inverses of the given rules. We ask it to generate inverse rules such that they contradict the given rules. We generate tests for both the rules from the OR and their inverses as shown in Figure 7. This approach helps us generate tests that cover intriguing test cases which might cause a given model to violate the OR.
@@ 3.3.3. Valid Tests
A test may not follow the IS and might create output that violates the rules described in the OR and PUT. Such test cases are not the most valuable, as the prompt developer knows that some of those inputs can be pre-filtered before they are passed to the prompt. However, these test cases might be useful for testing any input validation pipeline that processes the input or for identifying gaps in the assumed input domain. The most valuable tests follow the IS as they are actionable and require fixing, so we use the extracted IS to guide PromptPex in generating them. Even in scenarios where the prompt-based system is exposed to raw inputs, the same IS can validate inputs and disregard any input that does not meet the input specification.
@@
The test generator in PromptPex uses gpt-5_2025-08-07 and takes a rule (which can also be an inverse rule), IS and PUT as input. It generates a test along with the reasoning of why the model thinks this particular test during execution will comply with the rule it was generated for. We ask the test generator to start with first understanding: What is an input? What are the different components of a valid input? What are the syntax and semantics related constraints from IS? In the prompt, we direct the model to consider these issues during test generation. Each test must be generated such that the expected output or behavior demonstrates adherence to the given rules for a range of scenarios.
3.4. Test Evaluation: Compliance vs Correctness
@@
The generated test cases can be executed on multiple models, helping developers analyze prompt behavior across diverse inputs. As mentioned, because we lack the correct output for each test input, our automated strategy for test evaluation is an LLM-based determination of test compliance. Compliance is determined by our LLM-as-a-judge (Zheng et al., 2023) which, given the original prompt and the test output, determines whether the output complies with the requirements described in the prompt. As a result, our evaluation explores only a partial understanding of the model output and must be augmented with additional testing. However, as demonstrated in traditional software engineering, testing the assertions over the output can also be very valuable (Burnett et al., 2003; Clarke and Rosenblum, 2006; Taromirad and Runeson, 2024). After the tests are run, the results of our validator can assist the prompt developer in updating the prompt to correctly handle tests with invalid results.
@@
To implement LLM-as-a-judge, we use o4-mini_2025-04-16 and provide it the output generated by executing a test and the PUT. To allow a fair comparison with a test generator that uses PromptPex, we ensured that no artifacts specific to PromptPex are used in the validation and that the test output validator is generic, applicable to any prompt and test output regardless of the input, the method used to generate the test, or the model used to run the test.
This validator is used to evaluate the outputs of tests generated by both the baseline and PromptPex. We enforce the following properties for this output validator: 1) Keep the evaluation independent of the input, as it will not be provided. 2) Must not speculate, infer, or make any assumptions during the evaluation. 3) Always check for compliance, not correctness. The validator generates a boolean result (compliant/non-compliant) and explanation for the decision.
@@ 4. Evaluation
  1. (Q1)
    Do specifications generate better tests?
  2. (Q2)
    How useful are rules for generating better tests?
  3. (Q3)
    Does an input specification help generate valid tests?
  4. (Q4)
    Can automatically generated tests help determine if a model is suitable for a prompt?
4.1. Evaluation Setup
4.1.1. Baseline
In our evaluation, we used a zero-shot LLM-based test generator as the baseline. We use gpt-5_2025-08-07 to generate tests for both the baseline and PromptPex. We instruct the model to develop multiple test cases by inferring the functional specification and input for the given prompt. We ask it to begin with understanding possible inputs and its components, what the syntax and semantics related constraints are, and possible input scenarios. We enforce the following properties while generating the baseline tests: 1) The test cases must be designed to validate whether the output properly adheres to description, 2) A good test must always be a valid input meeting the requirements mentioned in the description, 3) The test cases must be diverse and distinct, 4) Each test case must be crafted to rigorously assess whether the output meets the stipulated behavior based on the provided prompt, 5) The input scenarios used for creating the tests must be valid, realistic, and fully comply with the given description, 6) Generate test cases to broadly cover a range of scenarios, including boundary cases, typical cases, and edge cases, to thoroughly evaluate the software’s adherence to the description, 7) Each test case should adhere to principles of good software testing practices, emphasizing coverage, specificity, and independence, 8) Focus on creating diverse test cases that effectively challenge the prompt’s capabilities by critically assessing potential weaknesses in the handling of inputs by the prompt.
We chose this as our baseline because LLMs are capable of interpreting the prompt and generating test cases for them. LLMs are already useful in generating unit tests for traditional software (Schäfer et al., 2024; Siddiq et al., 2024; Tang et al., 2024; Xie et al., 2023; Yuan et al., 2023, 2024; Yang et al., 2024). We ensured that the prompt used to generate the baseline tests is robust and underwent multiple revisions to enhance its ability to generate effective tests. To provide a fair comparison with PromptPex, we explicitly covered all requirements that are implicitly enforced by PromptPex, such as generating valid, challenging, and diverse test cases that comprehensively cover the prompt and help identify any flaws. Although there will always be opportunities to improve the baseline prompt, the same holds for the prompts used in PromptPex. We have refined both approaches sufficiently so that they can be effectively compared.
4.1.2. Metrics
@@
% non-compliance We define effective tests as those that expose limitations in the prompt, which means that they result in more failures. We consider non-compliance with the prompt as the metric for test quality. This includes any violations of the rules and constraints described in the prompt. This approach is similar to checking all the assertions over the output generated by a traditional program.
@@
We used LLM-as-a-judge to evaluate whether the output generated by the prompt violates any rules or constraints within the prompt as described in Section 3.4. We could have also developed a validator that does not use the PUT but instead relies on the OR, as they precisely represent the constraints over the output. However, since the OR is part of PromptPex and was not provided to the baseline test generator, we refrained from using it to maintain a fair and equal comparison. We run the tests generated by PromptPex and the baseline on the following models: gpt-oss, gemma2:9b, qwen2.5:3b, llama3.2:1b. We compare the non-compliance of the output from the generated tests. In our evaluation, we consider more test non-compliance as better since it indicates that the tests are more effective at identifying gaps in the prompt’s constraints.
Test Validity Valid tests follow the input spec (IS) defined by the prompt-under-test (PUT). We use test validity as a metric for PromptPex’s test quality. Valid tests are important as they represent likely inputs as compared with invalid inputs that may result in non-compliance but will also may be less likely to occur in actual use. We use an LLM-as-a-judge (o4-mini_2025-04-16) comparing the test against the IS to determine if the generated test is valid.
@@
Groundedness and Spec Agreement We also evaluate the generated OR as these are directly used to generate tests by PromptPex and have impact on the quality of the generated tests. A rule in the OR is considered grounded if it is present in the PUT. We do not want to generate tests for the rules which are not present in the PUT. We consider a higher groundedness of the rules as an indirect metric for test quality. To determine if the generated rules are grounded in the original PUT, we used LLM-as-a-judge and ask it to confirm rule groundedness.
@@
Figure 8. Spec agreement estimation for PromptPex tests. The specifications are extracted from the prompt and a spec prompt is created by appending the output specification to the task specification. Spec agreement is considered high when the spec prompt and original prompt behave similarly.
@@
Another goal for generating OR is to ensure that all important constraints mentioned in the PUT are captured in the OR. We refer to this as spec agreement. To compute this, we extract a description of how the output must be computed from the PUT, the task specification. We append the task specification with our extracted OR to derive a spec prompt which in theory should capture all the same constraints as the PUT. We do not use the IS in the spec prompt because PromptPex generates tests for each rule in the OR. To estimate how closely the extracted specification represents the original PUT, we compare the spec prompt with the original PUT. We feed tests generated from our baseline test generator to both the spec prompt and the PUT to compare their behaviors as shown in Figure 8. We use the baseline tests instead of the PromptPex tests so that they are not influenced by our process of generating them from the OR. The spec prompt and the original PUT must have similar behavior for the generated OR to have high spec agreement. We use cosine similarity of the non-compliance percentage of spec prompt and the original PUT to derive a spec agreement score.
4.1.3. Benchmarks